OpenStudy (anonymous):

Find the zeros of the function algebraically f(x)=9x^2-3x-2 Please explain step by step

OpenStudy (anonymous):

Step 1. Use quadratic formula. Step 2. Go get a snack.

OpenStudy (anonymous):

(You can omit step 2 if you want)

OpenStudy (anonymous):

Lol

OpenStudy (anonymous):

Could also be factored . . .

OpenStudy (anonymous):

factored?

OpenStudy (anonymous):

that's what I've been doing. I just don't know what to do with the 9x^2

OpenStudy (anonymous):

Yes like (2x+1)(x-3) = 2x^2-5x-3, (2x+1)(x-3) is the factored form of 2x^2-5x-3.

OpenStudy (anonymous):

okay, so that would make the 9x^2 go to (3x ) (3x ) ?

OpenStudy (anonymous):

Most likely. Recall that when you multiply two binomials, (ax+b)(cx+d) = acx^2+adx+bcx+bd.

OpenStudy (anonymous):

In your case a*c=9, and b*d=-2. You need to find a,b,c, and d such that ac=9, bd=-2 and ad+bc=-3

OpenStudy (anonymous):

sooo, it would come out to (3x-2) (3x+1) right?

OpenStudy (anonymous):

That looks right, so now you set each of those factors =0 to find the roots.